OpenStudy (anonymous):

Determine whether the graph of y = x2 − 6x + 3 has a maximum or minimum point, then find the maximum or minimum value.

OpenStudy (anonymous):

well this one you would need to use a graphing calculator to determine them but it easy enough to say it has none, anything that is a variation of y=x^2 is a parabola which is only a u-shape so if anything it has 1 minimum but if i remember they regard parabolas as none.

OpenStudy (anonymous):

The multiple choice answers that they are giving me is Minimum; (-6, 3) Maximum; (-6, 3) Minimum; (3, -6) Maximum; (3, -6)

OpenStudy (anonymous):

hmmm ok then there is a minimum

OpenStudy (anonymous):

So how would i do this

OpenStudy (anonymous):

make a simple graph of it, using a x/y chart and 3 simple points or if you can figureo ut it's vertex because that would be the minimum

OpenStudy (anonymous):

find the vertex

OpenStudy (anonymous):

Ohhhhh ok

OpenStudy (anonymous):

first coordinate of the vertex of \(y=ax^2+bx+c\) is \(-\frac{b}{2a}\)

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

in your case it is \(-\frac{-6}{2\times 1}=3\)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

the second coordinate of the vertex is what you get for \(y\) when you replace \(x\) by \(3\)

OpenStudy (anonymous):

OHhhh ok

OpenStudy (anonymous):

Then i would have to plug in x

OpenStudy (anonymous):

Which is 3

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

which i guess you do not have to do in this case, because the only choice is \((3,-6)\)

OpenStudy (anonymous):

YEHHHH I GOT THAT ANSWER 8'D /dance

OpenStudy (anonymous):

pick C because it is always C

OpenStudy (anonymous):

It would be a minimum right awesome

OpenStudy (anonymous):

yes, a minimum

OpenStudy (anonymous):

If i have more questions i will ask you guyz.

OpenStudy (anonymous):

k